OpenStudy (anonymous):

Solve each of these systems using the substitution method. 2x + 3y = -4 x − 2y = 5 Answer x = 1, y = -2 x = -1, y = 2 x = 1, y = 2 x = -2, y = -1 Question 2 2x − y = -2 3y = 6 − 5x Answer x = 0, y = 2 x = 2, y = 0 x = 2, y = 2 x = 0, y = -2 Question 3 2x − y = 4 6x − 3y = 7 Answer y = 2x - 4 No solution x = 2, y = -4 x = 2, y = 5/3 Question 4 2x − y = 3 4x = 6 + 2y Answer Infinitely many solutions x = 2, y = 1 y = 2x - 3 x = 4, y = 5 sos please help!

OpenStudy (shinalcantara):

Using substitution method 2x + 3y = -4 ----equation 1 x - 2y = 5 ----equation 2 from equation 2 x - 2y = 5 x = 2y + 5 ----equation 2' substitute equation 2' to equation 1 2x + 3y = -4 2(2y+5) + 3y = -4 4y + 10 + 3y = -4 4y + 3y = -4 - 10 7y = -14 y = -2 substitute y=-2 to equation 2' x = 2y + 5 x = 2(-2) + 5 x = -4 + 5 x = 1

OpenStudy (shinalcantara):

using the same method, can you do the next system of equations?

OpenStudy (anonymous):

im doing a math quiz rn bcs im homeschooled can u just help solve to other ones and thats it please it would mean alot

OpenStudy (shinalcantara):

but did you understand it though? (-_-)!

OpenStudy (anonymous):

i understand little bits of it.

OpenStudy (shinalcantara):

okay i'm gonna solve it but ask me those things that confuses you or those that you don't understand. okay?

OpenStudy (anonymous):

yeah sure

OpenStudy (shinalcantara):

2x - y = -2 ----equation 1 3y = 6 - 5x ----equation 2 from equation 1 2x - y = -2 2x + 2 = y y = 2x + 2 -----equation 1' substitute equation 1' to equation 2 3y = 6 - 5x 3(2x + 2) = 6 - 5x 6x + 6 = 6 - 5x 6x + 5x = 6 - 6 11x = 0 x = 0 substitute x=0 to equation 1' y = 2x + 2 y = 2(0) + 2 y = 2

OpenStudy (shinalcantara):

2x - y = 4 ----equation 1 6x - 3y = 7 ----equation 2 from equation 1 2x - y = 4 2x - 4 = y y = 2x - 4 ----equation 1' substitute equation 1' to equation 2 6x - 3y = 7 6x - 3(2x - 4) = 7 6x - 6x + 12 = 7 6x - 6x = 7 -12 0 is not equal to -5 for this case, there's no solution. there could be no values of x and y that would be true to both equations

OpenStudy (shinalcantara):

2x - y = 3 ---equation 1 4x = 6 + 2y ----equation 2 from equation 1 2x - y = 3 2x - 3 = y y = 2x - 3 ---equation 1' substitute equation 1' to equation 2 4x = 6 + 2y 4x = 6 + 2(2x - 3) 4x = 6 + 4x - 6 4x - 4x = 6 - 6 0=0 in this case where the variable becomes zero as the constant is zero, then we can say that the two equations are equal that whatever values of x and y you substitute, will be true to the two equations. the system does have an infinite solution

OpenStudy (anonymous):

but thats not in the choices o.o